OpenStudy (anonymous):

using the factoring process what constant is added to the right side of the quadratic equation. 3x^2+4x+____=1+_____ in order to solve by completing the square. 2 2/3 4/3 or 16 using the factoring process what constant is added to the right side of the quadratic equation. 2x^2-9x+____=-9+____ in order to solve by completing the square. 81/8 81/16 81/4 or 9 using the factoring process what constant is added to the right side of the quadratic equation. 2x^2+8x+_____=15+____ in order to solve by completing the square. 4 16 2 or 8

OpenStudy (anonymous):

sorry its kinda long

OpenStudy (dtan5457):

lol it's the same thing use (b/2)^2 for EVEERYTHHING on these questions.

OpenStudy (anonymous):

see the thing is i dont know how to work these problems out because i missed school for 6 weeks because i was in the hospital

OpenStudy (dtan5457):

i'm sorry to hear that but we just did these like 3 times :/

OpenStudy (dtan5457):

do you know what the b values are here ?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

3x^2+4x+____=1+_____ 3(x^2+(4/3)x+____)=1+3(_____)

OpenStudy (anonymous):

how do i figure that out

OpenStudy (tkhunny):

That was just factoring of the leading coefficient, 3.

OpenStudy (dtan5457):

ax^2+bx+c=0

OpenStudy (dtan5457):

they are all in that form

OpenStudy (anonymous):

im sorry i really dont know

OpenStudy (dtan5457):

@DanJS maybe your a better explainer than me

OpenStudy (anonymous):

thanks for trying to help im so sorry

OpenStudy (dtan5457):

it's ok. im just hoping someone can make this go through your mind better. i can't teach

OpenStudy (anonymous):

its just fine

OpenStudy (tkhunny):

We need this principle \((a+b)^{2} = a^{2} + 2ab + b^{2}\) We have \(x^{2} + (4/3)x + Something\) How do we decode that?

OpenStudy (anonymous):

i have no idea i havent learned any f this i was in the hospital and out of scool for 6 weeks

OpenStudy (tkhunny):

Okay, it requires thinking. Look at that "need" thing I wrote above. Ponder its properties. The first term is a perfect square. The third term is a perfect square. Can you build the second term, if you know only the 1st and 3rd terms? \(x^{2} + 2nd\;Term + 16\) What is the 2nd Term?

OpenStudy (anonymous):

I DO NOT KNOW

OpenStudy (tkhunny):

Not good enough. Think about it. Reread both of the last posts and answer the question.

OpenStudy (anonymous):

dont tell me its not good enough if i didnt learn it then i didnt learn it o there for i dont know it so no your answer isnt good enough you incompetent pig

OpenStudy (tkhunny):

I will tell you what I believe you need. Throwing your hands int he air is not good enough, by definition. Are you going to think about it or not? If you want to learn it and pass the exam, this requires thinking. The first term is a perfect square, \(x^{2}\). It is a square of what? The third term is a perfect square, \(16\). It is a square of what?

OpenStudy (anonymous):

wow you really dont get it maybe you should explain things a little better before you criticize anyone else about not understanding something learn to actually help then be an a hole stop trying to get past my level you wont ever reach it boo boo

OpenStudy (tkhunny):

Not criticizing. I'm giving you the facts and I am showing you how to think about it. If you want to control the engagement, you should consider paying for it. Since you are here for free, you should listen, rather than give orders. Answer my questions and you will learn. Go!

OpenStudy (dtan5457):

@boobear74 if you listen i guarantee this topic will be a lot easier

OpenStudy (dtan5457):

please listen.

OpenStudy (anonymous):

its hard to do something if i dont know what the dude is talking about at all so please do tell ,e how im supposed to listen

OpenStudy (dtan5457):

you can ask questions about what he said. but don't call him a pig or what not..

OpenStudy (dtan5457):

you need to ask a real question, not "i don't know"

OpenStudy (anonymous):

i did nothing was solved

OpenStudy (dtan5457):

no...you simply said "I don't know", but which part exactly?

OpenStudy (tkhunny):

The difficulty, here, is that you ASSUME you don't know what I'm talking about. In fact, you DO know. Let's try one question again. I think you CAN answer it. The third term is a perfect square, 16. It is a square of what?

OpenStudy (dtan5457):

@boobear74 Do you still want help??